OpenStudy (anonymous):

x+8y=7 (1) x=3-8y (2) solvefollowong the system of equations

OpenStudy (anonymous):

solve by elemate?

OpenStudy (anonymous):

rewite slope intercept x=3-8y ,x+8y =3 -x-8y =-3 x+8y =3 they are paralell

OpenStudy (anonymous):

does this mean there is no solution

OpenStudy (anonymous):

no intercept yes

OpenStudy (anonymous):

T Y

OpenStudy (anonymous):

yw do you know word problems?

OpenStudy (anonymous):

what problems?

OpenStudy (anonymous):

i will repost it here

OpenStudy (anonymous):

I try the best

OpenStudy (anonymous):

the equation y=-1777x+27,153 can be used to predict a number y of deaths after 2000 that is x=0 corresponds to 2000, x=3 corresponds to 2003, x=4 corresponds to 2004 and so on predict the number of gun deaths in 2004 and 2007 what year will the number of deaths be 12,937?

OpenStudy (anonymous):

sorry I don't know

OpenStudy (anonymous):

did you see any of my other posts you knew im really stumped on a lot of this im still learning basic algebra and its killing me lol

OpenStudy (anonymous):

I will look

OpenStudy (anonymous):

2004_ g(4) = -1777*4+27153 = 20045 ---------------- 2007_____ g(7) = -1777*7+27153 = 14714 ---------------- the predicted number of gun deaths will be 11,160 in the year_2009 12,937 = -1777x + 27153 12,937- 27153 = -1777x -14216\-1777 =x x= 8years

OpenStudy (anonymous):

here your answer good cluck

OpenStudy (anonymous):

your awesome thanks

OpenStudy (anonymous):

I will look another